OpenStudy (anonymous):

Find the length of the sides of a square if the number of square meters in its area is 1.5 times the number of meters in its perimeter.

OpenStudy (emmyma):

Hi! Find the length of the sides of a square if the number of square meters in its area is 1.5 times the number of meters in its perimeter. Let x = side of the square x^2 = 1.5(4x) x^2 = 6x x^2 - 6x = 0 x(x-6) = 0 x = 6 meters Good luck! Make sure to give me a medal if I helped. :)

OpenStudy (anonymous):

I have another one... If the perimeter of a square is 94m, find its area.

OpenStudy (emmyma):

Ask this in a different question so you can give me a medal! Tag me in it :)

OpenStudy (anonymous):

How do I tag u in it?

OpenStudy (emmyma):

@ emmyma without the space. :)